4. The Zaf Radiator Company uses a normal-costing system with a single manufacturing overhead cost pool and machine-hours asapliprint/accounting balances of Work-in-Process Control, Finished Goods Control, and Cost of Goods Sold. Begin by selecting(1) O Proportion (decimal) (2) (Over-) Underallocated OH Adjustment Allocation rate Ending Balance Allocation rate Beginning
Answer #1

1. BUDGETED MFG OH RATE AND UNDER / OVER APPLIED OH

budgeted mfg oh A          5,100,000
budgeted machine hours B                85,000
rate C= A/B                        60
Actual machine hours D                90,000
allocated overhead E = C X D          5,400,000
Actual mfg OH F          5,200,000
Over / (under) applied G = E-F              200,000

2.

2.a Balance before dispose of over applied OH overapplied OH Bal. after disposal of Over applied OH
cost of good sold                                  8,750,000 200000                        8,550,000
Finished Good control                                  3,125,000 0                        3,125,000
WIP Control                                     625,000 0                           625,000
                              12,500,000                     12,300,000

2B

2.b Disposal of over applied oh based on ending balance
Year end balance % % Over applied OH prop. Net balance
A B C D E F= D x E G= B- F
cost of good sold 8,750,000 8750000/12500000 0.700              200,000          140,000      8,610,000
Finished Good control 3,125,000 3125000/12500000 0.250              200,000            50,000      3,075,000
WIP Control 625,000 625000/12500000 0.050              200,000            10,000          615,000
12,500,000 1.000          200,000    12,300,000

2C

2.b Disposal of over applied oh based on overhead allocation rate as computed in first requirement
Year end balance Machine Hours Rate allocated overhead prop. prop. Over applied OH prop. Net balance
A B C D E = C X D F G H I= G X H J = B- I
cost of good sold                  8,750,000                63,000            60                      3,780,000 3780000/5400000              0.700                   200,000            140,000          8,610,000
Finished Good control                  3,125,000                18,000            60                      1,080,000 1080000/5400000              0.200                   200,000               40,000          3,085,000
WIP Control                      625,000                  9,000            60                         540,000 540000/5400000              0.100                   200,000               20,000              605,000
               12,500,000                      5,400,000              1.000            200,000        12,300,000

3

  • WHICH METHOD DO YOU PREFER ??
  • Alternative (c)  Disposal of over applied oh based on overhead allocation rate as computed in first requirement is theoretically peferred over (a) Disposal of over applied oh based by writting off to COGS and (b) Disposal of over applied oh based on ending balance as this method use actual machin hours for disposing over applied OH.
  • This alternative yeilds the same ending balance of WIP, Finished good and COGS had the actual idirect cost rate been used.